OpenStudy (anonymous):

if z=7-i/3-4i then find the value of z^14

OpenStudy (ash2326):

z=7-13i/3 |z|=root(49+169/9) |z|=root(610/9) no we'll find the argument tan theta= (-13/3)/7 theta= -31.7 degrees or theta=-0.552 radians so z=|z|e^(i*(-0.552)) z^14= (610/9)^7e^(i*(-0.552*7) z^14=(610/9)^7 e^(i(-3.85))